The mass of a molecule of a certain chemical compound is 2.45*10 to the power of -22 g. If each molecule is made up of two carbon atoms and six oxygen atoms and the mass of one oxygen atom is 1/3 that of a carbon atom, find the mass of an oxygen atom.
The answer is 2.04 * 10 to the power of -23 g. I do not understand this in terms of showing working.

Answer by edjones(7569)

Let c=carbon & g=oxygen
2c+6/3g=2.45*10^-22
2c+2g=2.45*10^-22 there are equal masses of oxygen and carbon in the compound.
0.5*2.45*10^-22
=1.225*10^-22 mass of oxygen in the compound.
(1/6)*1.225*10^-22 There were 6 oxygen atoms in the compound
=.2041667...*10^-22
=2.04...*10^-23 Mass of oxygen.
.
Ed